# LLM4SCREENLIT: Recommendations on Assessing the Performance of Large Language Models for Screening Literature in Systematic Reviews

### v1

@@ Abstract
Context: Large language models (LLMs) are released faster than users’ ability to evaluate them rigorously. When LLMs underpin research, such as identifying relevant literature for systematic reviews (SRs), robust empirical assessment is essential.
Objective: We identify and discuss key challenges in assessing LLM performance for selecting relevant literature, identify good (evaluation) practices, and propose recommendations.
Method: Using a recent large-scale study as an example, we identify problems with the use of traditional metrics for assessing the performance of Gen-AI tools for identifying relevant literature in SRs. We analyzed 27 additional papers investigating this issue, extracted the performance metrics, and found both good practices and widespread problems, especially with the use and reporting of performance metrics for SR screening.
Results: Major weaknesses included: i) a failure to use metrics that are robust to imbalanced data and do not directly indicate whether results are better than chance, e.g., the use of Accuracy, ii) a failure to consider the impact of lost evidence when making claims concerning workload savings, and iii) pervasive failure to report the full confusion matrix (or performance metrics from which it can be reconstructed) which is essential for future meta-analyses. On the positive side, we extract good (evaluation) practices on which our recommendations for researchers and practitioners, as well as policymakers, are built.
Conclusions: SR screening evaluations should prioritize lost evidence/recall alongside chance-anchored and cost-sensitive Weighted MCC (WMCC) metric, report complete confusion matrices, treat unclassifiable outputs as referred-back positives for assessment, adopt leakage-aware designs with non-LLM baselines and open artifacts, and ground conclusions in cost–benefit analysis where FNs carry higher penalties than FPs.
@@
[inst1]organization=Wrocław University of Science and Technology, addressline=Wyb. Wyspianskiego 27, city=Wrocław, postcode=50-370, country=Poland
@@ 1 Introduction
Large language models (LLMs) are increasingly employed to automate the challenging task of paper screening in systematic reviews (SRs), promising substantial reductions in human workload and faster evidence synthesis in software engineering [1, 2, 3] and beyond [4, 5, 6, 7, 8, 9, 10, 11]. To quantify their effectiveness, standard confusion-matrix metrics, e.g., accuracy, precision, recall, specificity, and F1-score, are adopted by comparing model decisions (include/exclude) against human reference labels. Although these metrics offer a familiar evaluation framework, their uncritical application can yield misleading conclusions. In particular, we argue it is essential to consider the features of the problem domain and which metrics best address them. This contrasts with an evaluation using all the ‘usual’ metrics in the hope that embedded within the results will be useful insights.
@@
We argue that there are four key features relating to screening papers for SRs. First, the data will tend to be extremely imbalanced so that the negative class (i.e, papers that are irrelevant to the SR) will considerably outnumber the positive class (i.e., relevant papers). Second, the costs of misclassifications are not equal. A relevant paper wrongly excluded, i.e., a false negative (FN), will likely have a far greater impact upon the quality of the SR than an irrelevant paper that passes the screening, i.e., a false positive (FP), and then wastes human effort subsequently rectifying the situation. Third, resources are limited, so we are concerned about the overall costs and benefits of deploying different screening tools. Fourth, we would like to be reassured that sophisticated, yet essentially black-box methods such as LLM screening tools are actually doing better than guessing.
Delgado-Chaves et al. [4] (hereafter referred to as DC+) recently evaluated 18 LLMs for screening studies in three SRs, comparing selections with human reviewers and using confusion matrix metrics. They also emphasize comparing LLMs with more traditional machine learning methods, specifically the random forest method. While DC+ is a key step in assessing LLMs for SR screening, methodological issues (using simple counts of correctly classified papers—referred to as Accuracy in classification studies—as performance metrics, reporting biased metrics, and excluding unclassifiable papers from performance assessment) obscure its outcomes.
@@
We also report the results of reviewing 27 other papers that studied the use of LLMs to screen literature for SLRs. The papers were obtained from the primary studies included in two systematic reviews ([12] and [13]), together with papers we found from informal searches at the beginning of our investigation. We investigated these papers for three purposes: i) to confirm that the problems we observed in the DC+ paper are not unique to that paper, ii) to assess whether any other problems exist, and iii) to investigate whether there are additional good practices to recommend.
@@ 2.1 The Fallacy of Correctness
This statement is misleading because the reviews SR-I and SR-II are highly imbalanced (many irrelevant vs. few relevant studies), meaning rejecting all studies would still achieve high scores for correctness and percentage correctness. For example, gemma:7b in Table 1 scores 96.17% for Accuracy but found none of the relevant studies (TP=0). Similarly, mistral-nemo:12b scored 96.11% for Accuracy found none of the relevant studies, and identified 3 FPs (i.e., irrelevant studies) as relevant. In contrast, two models that at least identified some of the relevant papers (TP>>0) had lower Accuracy values. This means that if we optimise on the Accuracy metric, we would select worse, or in some cases, completely ineffectual LLMs that failed to detect any relevant primary studies.
@@
Table 1: Performance metrics for four of the LLMs used in SR-I, revealing problems with using Accuracy and other non-chance adjusted metrics, and ignoring relative costs. NB The asterisks ‘*’ denote the ‘best’ LLM by metric, i.e., row-wise. The double asterisks ‘**’ denotes that we used weight $w=10$ to calculate WMCC.
@@ 2.2 The Critical Importance of Lost Evidence
In SRs, falsely rejecting relevant studies (FNs) loses evidence, potentially irretrievably. Recall, also referred to as Sensitivity (the proportion of relevant studies identified), and Lost Evidence (1-Recall) are therefore fundamental performance metrics. DC+ Figure 1 reveals all reviews—even the balanced SR-III—had problematic Lost Evidence scores. Figure˜1 shows that Lost Evidence was serious across all models. Across all SRs, Lost Evidence ranged from 14% (best-case in SR-II) to 100% (worst case in SR-I), with 46 out of 54 LLM classifications missing more than 50% of positive papers. No model performed well on all datasets. The only consistency was llama3.2:3b delivered extremely poor predictions on all three data sets, likely because it had the fewest parameters.
Figure 1: Lost Evidence per Model for three SRs (the median of Lost Evidence is presented as a point) and the min/max show the extremes)
Unlike FNs, FPs (irrelevant studies incorrectly included) only waste effort in subsequent screening. The dominant risk to SR validity comes from missing papers that should be included. However, allowing an extremely large number of FPs, in order to minimize FNs, would mean that there was little value in using the LLM classification. This can be modelled as the FN/FP cost ratio, requiring a subjective assessment of the cost of missing evidence versus the cost saving involved in not processing irrelevant studies. The cost ratio will vary by domain (healthcare SRs may value lost evidence more highly than software engineering SRs) and review type (systematic scoping reviews tolerate missing studies better than formal SRs).
@@ 2.3 Reporting Biased and Unsuitable Performance Metrics
Of the metrics DC+ deploy, only the Matthews Correlation Coefficient (MCC) [17] is unbiased because it considers all four elements of the confusion matrix without any bias towards TPs or TNs. MCC is an application of the Pearson correlation coefficient, ranging from -1 to 1, with values near zero indicating chance-level performance. Like any correlation coefficient, MCC remains reasonably robust to imbalanced datasets [17, 18]. The problem is that MCC does not address the differential costs of FPs and FNs. However, it is possible to construct an appropriate weighted measure of MCC (WMC), which we discuss in Section˜3.3.
In addition, reporting multiple performance metrics is inherently misleading because all these metrics are derived from the same four confusion matrix elements, and the elements of the confusion matrix are not themselves independent. This is clear because we only require limited information about the overall classification process in order to generate all four elements. For example, if we know the number of negative papers (N), the number of positive papers (P) as defined by the baseline (gold standard) classification process, together with the number of papers classified as negative by the LLM (n) and the number of those n papers that were TNs, then we can to construct the remaining three elements of the confusion matrix because:
@@
Thus, a large number of related performance metrics are unhelpful because they are functionally correlated in ways that are more often more difficult to understand than the basic confusion matrix elements.
@@ 2.4 Dropping Unclassified Papers
2.5 Good Practices
Although we have criticised its choice of performance metrics, the DC+ paper adopts two extremely useful good (evaluation) practices:
(P1) Reporting full confusion matrices: The paper provides the full confusion matrices for each LLM and SR in publicly accessible supplementary materials. Access to the full confusion matrices means that other researchers and meta-analysts can easily construct any performance metric of interest in their own context, in particular, the unbiased MCC metric or the weighted MCC metric.
(P2) Comparing with non-LLM baselines: The paper also assesses the performance results obtained when using a more traditional classification technique (i.e., the random forest method). This is important because researchers wanting to improve the efficiency of their SR process are not concerned only about the relative effectiveness of different LLMs, but also need to know whether LLMs outperform other techniques that do not share the problems associated with LLM use, such as the risk of hallucinations.
@@ 3 Current LLM Literature Screening Evaluation Practices
To assess whether the performance metric problems in the DC+ paper were representative of current research practice, we also reviewed another 27 papers addressing literature screening, see LABEL:tab:SummaryOfScreeningPapers.
@@
Guo-2024 [30] Acc; F1; Kappa; PABAK; RecExc; RecInc Yes Both J No
@@
Khraisha-2024 [6] Acc; Kappa; PABAK; Rec; Spec; Weighted Kappa No Kim J No
@@
Wang-2024 [10] BalAcc; F3; Prec; Rec; No Both C No
@@
    The papers were assembled from three different sources, which are shown in the columns labelled Origin. Papers were obtained from two systematic reviews of papers reporting empirical studies of LLM support for literature selection: Kim et al. [12] included 14 papers reporting 33 separate studies in their meta-analysis of the performance metrics F1, precision and Recall/Sensitivity, and Sander et al. [13] included 11 papers reporting 13 separate studies in their systematic review of Recall/Sensitivity and workload reduction. Five papers were included in both studies (labelled as Both in the Origin column). In addition, we identified 10 other papers (including DC+) from our own informal searches. These are labelled A (for Authors) in the Origin column. Only two of the papers were also identified by Kim et al. [12] or Sander et al. [13].
@@
    The column labelled Performance Metrics, identifies the metrics reported in each paper, excluding simple confusion metric counts. The performance metrics referred to by shortened labels are Acc (Accuracy or Correctness), BalAcc (Balanced Accuracy), Rec (Recall or Sensitivity), Spec (specificity), Prec (Precision), NegPred (Negative Prediction), RecInc (Sensitivity Included), RecExc (Sensitivity Excluded, Pos (Positive), Neg (Negative), Null (Number of null or otherwise invalid outcomes), PropMissed (Proportion Missed). The term WS was used to refer to some form of work saved metric irrespective of the specific term used by the authors. In addition, Akinseloyin et al [22] investigated prioritising papers in terms of relevance, rather than classifying them. They used metrics appropriate to that task, but did not fully define them. Their Recall statistics and Work saved statistics were calculated relative to the percentage of prioritised papers evaluated.
@@
    The six papers that reported confusion metric counts or percentages are identified in the column labelled CM.
@@
  5. 5.
    The column labelled Other Baselines identifies the papers that compared LLM performance with other types of machine learning algorithms such as logistic regression.
@@
Figure 2: Distribution of evaluation metrics used across 27 papers analyzing Gen-AI tools for systematic review screening

### v2

@@ Abstract
Context: Large language models (LLMs) are increasingly used to screen literature for systematic reviews (SRs), but the standard confusion-matrix metrics used to evaluate them can mislead under the imbalanced, cost-asymmetric conditions of screening.
Objective: We develop and justify LLM4SCREENLIT—practical recommendations for researchers conducting LLM-screening evaluations and for editors and reviewers assessing such studies—differentiated by study type (retrospective benchmarking vs. deployment for a specific SR).
Method: Using Delgado-Chaves et al. (2025), an 18-LLM benchmark across three biomedical SRs, as motivating example, we reviewed 28 additional papers and extracted their reported metrics. We propose a Weighted Matthews Correlation Coefficient (WMCC) that integrates MCC’s chance-correction with asymmetric misclassification costs, and validated it on three software-engineering (SE) reanalyses (Felizardo et al. 2024; Syriani et al. 2024; Huotala et al. 2025), the largest covering 9 LLMs $\times$ 24 SE secondary studies (34,528 articles).
Results: Across the 29 papers, only 10% reported MCC, only 24% reported full confusion matrices, and none of the five papers claiming workload savings priced false-negative cost. In the largest SE reanalysis, MCC and WMCC disagree on the best LLM in 55% of evaluable studies; in the most striking 9,695-article SE study, the Accuracy-best LLM loses 63.3% of relevant evidence (Lost Evidence), the MCC-best 43.9%, but the WMCC-best only 5.8%. Sensitivity analysis (median crossover at $w\approx 2.7$, all <7<7) supports $w=10$ as a conservative default.
Conclusions: SR-screening evaluations should prioritize Lost Evidence and use cost-sensitive WMCC alongside MCC for ranking. Reporting must include the full confusion matrix and treat unclassifiable outputs as positives requiring human review. Designs should be leakage-aware, with non-LLM baselines when the study aims to inform SR practice and labels are available. Editors and reviewers should require these elements as routine. Extension to full-text screening and data extraction is principled but pending empirical validation.
@@
[inst1]organization=Wroclaw University of Science and Technology, addressline=Wyb. Wyspianskiego 27, city=Wroclaw, postcode=50-370, country=Poland
@@ 1 Introduction
Large language models (LLMs) are increasingly employed to automate the challenging task of paper screening in systematic reviews (SRs), [^1] Systematic reviews are also referred to as systematic literature reviews or SLRs in the software engineering literature. We use SR throughout this paper but retain SLR where cited studies use it as part of their own terminology (e.g., Felizardo et al.’s dataset names SLR1 and SLR2). promising substantial reductions in human workload and faster evidence synthesis in software engineering [1, 2, 3, 4, 5, 6] and beyond [7, 8, 9, 10, 11, 12, 13, 14]. To quantify their effectiveness, standard confusion-matrix metrics, e.g., accuracy, precision, recall, specificity, and F1-score, are adopted by comparing model decisions (include/exclude) against human reference labels. Although these metrics offer a familiar evaluation framework, their uncritical application can yield misleading conclusions. In particular, we argue it is essential to consider the features of the problem domain and which metrics best address them. This contrasts with an evaluation using all the ‘usual’ metrics in the hope that somehow useful insights might emerge.
@@
We argue that there are four key features relating to screening papers for SRs. First, the data will tend to be extremely imbalanced so that the negative class (i.e, papers that are irrelevant to the SR) will considerably outnumber the positive class (i.e., relevant papers). Second, the costs of misclassifications are unequal. A relevant paper wrongly excluded, i.e., a false negative (FN), will likely have a far greater impact upon the quality of the SR than an irrelevant paper that passes the screening, i.e., a false positive (FP), and then wastes human effort subsequently rectifying the situation. Third, resources are limited, so we are concerned about the overall costs and benefits of deploying different screening tools. Fourth, we would like to be reassured that sophisticated, yet essentially black-box methods such as LLM screening tools are actually doing better than guessing.
While the studies we review span multiple domains, including biomedicine, the methodological challenges we address—class imbalance, asymmetric misclassification costs, and the need for chance-anchored metrics—are fundamentally mathematical rather than domain-specific. These recommendations are directly relevant to the growing adoption of LLMs for SR screening in SE [1, 4, 3, 5, 6].
Delgado-Chaves et al. [7] (hereafter referred to as DC+) recently evaluated 18 LLMs for screening studies in three SRs covering physiotherapy, neurology, and digital health topics (hereafter referred to as SR-I, SR-II, and SR-III), comparing selections with human reviewers and using confusion matrix metrics. They also emphasize comparing LLMs with more traditional machine learning methods, specifically the random forest method. While DC+ represents a valuable and ambitious contribution to assessing LLMs for SR screening, we use it as an illustrative example to highlight methodological issues that are common across the fields—such as reporting Accuracy as a primary metric under class imbalance—because DC+ provides complete confusion matrices that enable detailed reanalysis. Note that despite being drawn from the biomedical domain, we have used DC+ as our motivating example because it, in some regards, represents good practice, is published in a prestigious and influential venue, and because many pioneering and important methodological advances have come originally from this domain and then have subsequently been adopted by fields such as software engineering.
@@
We also report the results of reviewing 28 other papers that studied the use of LLMs to screen literature for SRs. The papers were obtained from the primary studies included in two systematic reviews ([15] and [16]), together with papers we found from informal searches at the beginning of our investigation. We investigated these papers for three purposes: i) to confirm that the problems we observed in the DC+ paper are not unique to that paper, ii) to assess whether any other problems exist, and iii) to investigate whether there are additional good practices to recommend.
The remainder of the paper is organised as follows. First, in Section 2 we describe our motivating study DC+ and frame it in the context of confusion matrices and how these provide a useful abstraction. From here we proceed (in Section 3) to the current, wider literature on LLM-based screening of studies for SRs, including SE-specific worked examples that illustrate the impact of metric choice on SE evaluations (Section˜3.4). From these two sources, we then (Section 4) derive the LLM4SCREENLIT recommendations and implications, structured by study type (benchmarking vs. deployment), and conclude with some limitations of the study and suggestions for further work.
@@ 2.1 The Fallacy of Correctness
This statement is misleading because the reviews SR-I and SR-II are highly imbalanced (many irrelevant vs. few relevant studies), meaning rejecting all studies would still achieve high scores for correctness and percentage correctness. For example, gemma:7b in Table 1 scores 96.17% for Accuracy but found none of the relevant studies (TP=0). Similarly, mistral-nemo:12b scored 96.11% for Accuracy, found none of the relevant studies, and identified 3 FPs (i.e., irrelevant studies) as relevant. In contrast, two models that at least identified some of the relevant papers (TP>>0) had lower Accuracy values. This means that if we optimise on the Accuracy metric, we would select worse, or in some cases, completely ineffectual LLMs that failed to detect any relevant primary studies.
@@
Table 1: Performance metrics for four of the LLMs used in SR-I, revealing problems with using Accuracy and other non-chance adjusted metrics, and ignoring relative costs. NB The asterisks ‘*’ denote the ‘best’ LLM by metric, i.e., row-wise. The double asterisks ‘**’ denotes that we used a weight $w=10$ to calculate WMCC.
@@ 2.2 The Critical Importance of Lost Evidence
In SRs, falsely rejecting relevant studies (FNs) loses evidence, potentially irretrievably. Recall, also referred to as Sensitivity (the proportion of relevant studies identified), and Lost Evidence (1-Recall) are therefore fundamental performance metrics. DC+ Figure 1 reveals all reviews—even the balanced SR-III—had problematic Lost Evidence scores. Figure˜1 shows that Lost Evidence was problematic across all models, ranging from 14% (best-case in SR-II) to 100% (worst case in SR-I), with 46 out of 54 LLM classifications missing more than 50% of positive papers. No model performed well on all datasets. The only consistency was llama3.2:3b delivered extremely poor predictions on all three data sets, likely because it had the fewest parameters.
Figure 1: Lost Evidence per Model for three SRs (the median of Lost Evidence is presented as a point and the min/max show the extremes)
Unlike FNs, FPs (irrelevant studies incorrectly included) only waste effort in subsequent screening. The dominant risk to SR validity comes from missing papers that should be included. However, allowing an extremely large number of FPs, in order to minimize FNs, would mean that there was little value in using the LLM classification. This can be modelled as the FN/FP cost ratio, requiring a subjective assessment of the cost of missing evidence versus the cost saving involved in not processing irrelevant studies. The cost ratio will vary by domain and review type (e.g., systematic scoping reviews may tolerate missing studies better than formal SRs), and should be determined by stakeholders based on the specific context and consequences of missed evidence.
@@ 2.3 Reporting Biased and Unsuitable Performance Metrics
Of the metrics DC+ deploy, only the Matthews Correlation Coefficient (MCC) [20] is unbiased because it considers all four elements of the confusion matrix without any bias towards TPs or TNs. MCC is an application of the Pearson correlation coefficient, ranging from -1 to 1, with values near zero indicating chance-level performance. Like any correlation coefficient, MCC remains reasonably robust to imbalanced datasets [20, 21]. The problem is that MCC does not address the differential costs of FPs and FNs. However, it is possible to construct an appropriate weighted measure of MCC (WMCC), which we discuss in Section˜3.3.
In addition, reporting multiple performance metrics without adequate interpretive framing can be problematic because all these metrics are derived from the same four confusion matrix elements, and the elements of the confusion matrix are not themselves independent. Multiple metrics can be useful if properly contextualized, but the functional correlations between them are often more difficult to understand than the basic confusion matrix elements. This is clear because we only require limited information about the overall classification process in order to generate all four elements. For example, if we know the number of negative papers (N), the number of positive papers (P) as defined by the baseline (gold standard) classification process, together with the number of papers classified as negative by the LLM (n) and the number of those n papers that were TNs, then we can construct the remaining three elements of the confusion matrix because:
@@ 2.4 Dropping Unclassified Papers
2.5 Good Evaluation Practices
While DC+ illustrates the metric selection challenges discussed above, the paper also adopts two extremely useful good evaluation practices:
(P1) Reporting full confusion matrices: DC+ provides the full confusion matrices for each LLM and SR in publicly accessible supplementary materials. Access to the full confusion matrices means that other researchers and meta-analysts can easily construct any performance metric of interest in their own context, in particular, the unbiased MCC metric or the weighted MCC metric (see Section 3.3).
(P2) Comparing with non-LLM baselines: Non-LLM baselines complement LLM evaluation when the study goals and study design support them (typically SR updates with prior screening labels, or retrospective benchmarks that include labelled datasets). In particular, non-LLM baselines are often informative when the study aims to assess whether LLMs can support or integrate with existing SR practice. In contrast, for studies whose sole aim is to compare LLMs’ performance against each other, or to evaluate different prompting strategies, non-LLM baseline comparisons are unnecessary and P2 can be ignored.
@@ 3 Current LLM Literature Screening Evaluation Practices
To assess whether the performance metric problems in the DC+ paper were representative of current research practice, we also reviewed 28 additional papers on literature screening, including any supplementary materials (see LABEL:tab:SummaryOfScreeningPapers).
The papers were assembled from three different sources, which are shown in the columns labelled Origin. Papers were obtained from two systematic reviews of papers reporting empirical studies of LLM support for literature selection: Kim et al. [15] and Sandner et al. [16]. Kim et al. undertook a meta-analysis of the performance metrics F1, Precision, and Recall/Sensitivity based on 14 relevant papers (including one SE paper [2]) that together reported 33 separate results. Sandner et al. undertook an SR using the performance metrics Recall/Sensitivity and workload reduction. They identified 11 relevant papers (none being SE related) reporting 13 separate results. Five papers were included in both SRs (labelled as Both in the Origin column). In addition, we identified 14 relevant papers (including DC+) from our own informal searches. These papers are labelled A (for Authors) in the Origin column of LABEL:tab:SummaryOfScreeningPapers. Five of the papers we found were also identified by Kim et al. [15] or Sandner et al. [16]. In addition to [1], we initially identified another five SE-related papers (i.e., [4][5], [6], [3] and [2]), bringing the total number of papers assessed to 29.
@@
Guo-2024 [33] Acc; F1; Kappa; PABAK; RecExc; RecInc Yes Both;A J No
@@
Khraisha-2024 [9] Acc; Kappa; PABAK; Rec; Spec; Weighted Kappa No Both;A J No
@@
Wang-2024 [13] BalAcc; F3; Prec; Rec; No Both;A C No
@@
    The column labelled Performance Metrics, identifies the metrics reported in each paper, excluding simple confusion metric counts. The performance metrics referred to by shortened labels are Acc (Accuracy or Correctness), BalAcc (Balanced Accuracy), Rec (Recall or Sensitivity), Spec (specificity), Prec (Precision), NegPred (Negative Prediction), RecInc (Sensitivity Included), RecExc (Sensitivity Excluded, Pos (Positive), RecAT% (Recall at % checked from an ordered list), Neg (Negative), Null (Number of null or otherwise invalid outcomes), PropMissed (Proportion Missed). The term WS was used to refer to some form of work saved metric irrespective of the specific term used by the authors. In addition, Akinseloyin et al [25] investigated prioritising papers in terms of relevance, rather than classifying them. They used metrics appropriate to that task, but did not fully define them. Their Recall statistics and Work saved statistics were calculated relative to the percentage of prioritised papers evaluated.
@@
    The seven papers that reported confusion metric counts or percentages (in the paper itself or in supplementary material from which counts can be reconstructed) are identified in the column labelled CM.
@@
    The column labelled Other Baselines identifies the papers that compared LLM performance with other types of machine learning algorithms, such as logistic regression.
@@
Figure 2: Distribution of evaluation metrics used across 29 papers analyzing Gen-AI tools for systematic review screening